OpenStudy (anonymous):

There are two different compounds of sulfur and fluorine. In SF6, the mass of fluorine per gram of sulfur is 3.55 g F/g S. In the other compound, SFX, the mass of fluorine per gram of sulfur is 1.18 g F/g S. What is the value of X for the second compound? Express your answer as an integer.

OpenStudy (anonymous):

in SFx you have one Sulphur its molecular weight=32gm so,weight of flourine=32*1.18=38gm divide it by molecular weight of flourine to get no of flourines i.ex

OpenStudy (anonymous):

x=38/19=2